I need help solving this equation using the substitution method. Can you please answer it as detailed as possible.
2x - 3y = 0
-4x + 3y = -1

Answer by mananth(16946)

2 x -3 y = 0 .............1
3 y = -1 + 4 x ---------2
/ 3 =
y = -0.33 + 1.33 x
Substitute y in (1)
2x-3 ( -0.33 + 1.33 x ) = 0
2 x + 1 + -4 x = 0
2 x -4 x = 0 -1
-2 x = -1
/ -2
x= 0.5
Plug the value of x in (1)
2 x -3 y = 0
2 * 0.5 -3 y = 0
1 -3 y = 0
-3 y = -1
/ -3
y= 0.33
m.ananth@hotmail.ca